OpenStudy (anonymous):

(4x^2/9)+(9/25)=(4x)/(5)

OpenStudy (anonymous):

I know until the \[(\frac{ 2x }{ 3 }-\frac{ 3 }{ 5 })^{2}\]

OpenStudy (anonymous):

arrives.

OpenStudy (anonymous):

\[(\frac{2x}{3}-\frac{3}{5})^2=0\] \[\frac{2x}{3}=\frac{3}{5}\]

OpenStudy (anonymous):

\[x=\frac{9}{10 }\]

OpenStudy (anonymous):

Thank you.